Solve this system of equations using substitution. Put your answer in ordered pair form, (x, y).
5x + y = 1
3x + 2y = 2

Answer by checkley79(3341)

5x + y = 1 OR Y=1-5X
3x + 2y = 2
3X+2(1-5X)=2
3X+2-10X=2
-7X=2-2
-7X=0
X=0 ANS.
3*0+2Y=2
2Y=2
Y=2/2
Y=1 ANS.
PROOF:
5*0+1=1
1=1